OpenStudy (anonymous):

solve the following system of equations. 2x = 7+4y 4y + 3x = -2

OpenStudy (anonymous):

2x - 4y = 7 3x + 4y = -2

OpenStudy (anonymous):

x=7/2+2y 4y+2(7/2+2y)=-2 4y+7+4y=-2 8y=-2-7 8y=-9 8y/8=-9/8 y=-9/8

OpenStudy (anonymous):

I'd start with Velvet's approach. Rearrange the first equation so that it is in standard form (ax + by = c) by subtracting 4y from each side. 2x - 4y = 7 3x + 4y = -2 Now, we want to combine them in a way that gets rid of one variable -- in this case, it's easy to get rid of y by adding the two equations together. (-4y + 4y = 0) 5x = 5 divide both sides by 5 to get x=1 plug this in to either original equation, to get y: 2(1) = 7+4y 2 = 7 + 4y -5=4y -5/4 = y check it by plugging out answers (1, -5/4) into the second equation: 3(1) + 4(-5/4) = -2 3 + (-20/4) = -2 3-5 = -2 -4y = 5

OpenStudy (anonymous):

Didn't realize there's a length limit -- to finish: y = -5/4 solution = (1, -5/4), which can be checked by plugging into the second equation. 3(1) + 4(-5/4) = -2, which it does